Using Gauss's theorem, calculate the flux of the vector field A vector=x³i^+x²zj^+yzk^ through the surface of a cube of side 2 units.


1
Expert's answer
2021-01-23T14:07:33-0500
"div {\vec{A}}=3x^2+0+y=3x^2+y\\
\Phi=\int_0^2\int_0^2\int_0^2(3x^2+y)dxdydz\\
\Phi=2\int_0^2\int_0^2(3x^2+y)dxdy\\
\Phi=2\int_0^2((2)3x^2+0.
5(2)^2)dx=40\\"